Three point charges are fixed in place in a right triangle. What is the electric force on the +1.5 HC charge due to the other two charges? +0.84 uC 7.9 cm \ T 0.3 cm .y -0.68 uC +1.5 uC
Answer #1

Please let me know if there is any mistake.

Here is the free body diagram,

0.103 0.079 45 0.06609

Here the angle is not 45 degree

sin theta = 7.9 / 10.3

theta = 50.084 degree

In x direction we have

Fx = - 9e9 * 0.68e-6 * 1.5e-6 / 0.066092 + 9e9 * 0.84e-6 * 1.5e-6 * cos 50.084 / 0.1032

Fx = - 2.1017 + 0.6858

Fx = -1.4158 N

In y direction

Fy = - 9e9 * 0.84e-6 * 1.5e-6 * sin 50.084 / 0.1032

Fy = - 0.8198 N

so,

F = sqrt ((-1.4158)2 + (-0.8198)2)

F = 1.636 N

direction would be

theta = 30 DEGREE below the negative x axis.